What quantity of 60 per cent acid solution must be mixed with a 20 percent solution to produce 720 mL of a 50 per cent solution?
Answer by jrfrunner(365)

Let X = amount of 60% solution needed
Let Y = amount of 20% solution needed
==
You are given information to form two equations and you have two unknowns
which are necessary conditions to solve the problem
equation 1 (amounts): X+Y =720ml
equation 2 (concentrations): 0.6*X+0.2*Y=0.5(720)
===
solve equation 1 for one of the variables (ie Y): Y=720-X
multiply equation 2 by 10 to make into whole numbers
6*X+2Y=5*720
substitute Y=720-X into previous equation
6*x+2*(720-X)=3600
6x+1440-2x=3600
4x=3600-1440=2160
x=540
===
y=720-x
y=720-540=180